Please help.
a boat can go 30 miles down stream 3/5th the time that it takes to return the same distance upstream. If the rate of the current is 3 miles per hour, what is the rate of the boat in still water?
a. 8mph
b. 10mph
c. 12mph
d. 16mph

Answer by mananth(16949)

speed of boat = xmph
current speed = 3 mph
..
upstream speed = x-3 mph
downstream speed = x+3 mph
...
distance = 30 miles
(3/5) *30/(x-3) = 30/(x+3)
18/(x-3)=30/(x+3)
18(x+3)=30(x-3)
18x+54=30x-90
30x-18x=90+54
12x=144
/12
x= 12 mph